For a 3 x 5 ANOVa with 9 subjects in each cell:

a. Find the value of each df component, and show that they sum to the df total

b. Find the critical value for F for each main effect and the interaction

Please show all work and how you found each main effect and the interaction.
Answer #1

(a) df of factor A = 3 - 1 = 2

df of factor B = 5 - 1 = 4

df of the interaction = 2*4 = 8

df of the error = 3*5*(9 - 1) = 120

df of the total = 2 + 4 + 8 +120 = 134

(b)

Critical value
Factor A 3.0717794
Factor B 2.4472365
Interaction 2.0164256

The formulas are:

Critical value
Factor A =FINV(0.05,2,120)
Factor B =FINV(0.05,4,120)
Interaction =FINV(0.05,8,120)

Please give me a thumbs-up if this helps you out. Thank you!